the locus of the mid pt. of a chord of the circle x2 +y2 =4,which subtend a right angle at the origin is:

Grade:12

1 Answers

Chetan Mandayam Nayakar
12 years ago

radius of circle=2

so length of chord=2√2

let O be origin and P be mid-point of chord

OP=sqrt(4-2)=√2

so locus is x2+y2=2